OpenStudy (anonymous):

Find the radius of a sphere with a surface area of 804 cm

OpenStudy (anonymous):

Multiple Chose: 9 cm 8 cm 64 cm 204 cm

OpenStudy (anonymous):

surface area of the sphere = \[4 \pi r^2\] \[804 = 4 \pi r^2 \rightarrow 804 = 4 \times \frac{22}{7} \times r^2\] \[r^2 = \frac{ 804 \times 7}{4 \times 22}=\frac{ 5628}{88}=63.95 \approx 64 cm\] \[r^2= 64 cm \rightarrow r= 8 cm\] is the radius o the sphere. @edengrace001

OpenStudy (anonymous):

Thank you soooo much! @dpasingh

OpenStudy (anonymous):

@edengrace001 Medal please..